If a sample of 25 canned goods is selected, what is the probability that the mean of the sample will be larger than 2210mg?


1
Expert's answer
2020-06-04T20:28:34-0400

Let "X=" the mean of the sample :"X\sim N(\mu=215, \sigma^2=15^2)." Then "Z=\dfrac{X-\mu}{\sigma}\sim N(0, 1)"

The probability that the mean of the sample will be larger than 220mg is


"P(X>220)=1-P(X\leq220)="

"=1-P(Z\leq {220-215\over 15})\approx1-P(Z\leq0.3333)\approx"

"\approx0.3694"